OpenStudy (anonymous):

Show that for any triangle (a^2+b^2+c^2)/(2abc)=(cos A/a)+(cos B/b)+(cos C/c)

OpenStudy (experimentx):

expand each square term on left using cosine rule. (a^2+b^2+c^2)/(2abc) = (b^2+c^2 - 2bc cosA + a^2+c^2 - 2 ac cos B +a^2+b^2 - 2 ab cos C)/2 = 2(a^2 + b^2 + c^2)/2abc -1/2( (cos A/a)+(cos B/b)+(cos C/c) ) let (a^2+b^2+c^2)/(2abc) = x then 1/2 x = x - 1/2((cos A/a)+(cos B/b)+(cos C/c)) 1/2 x = 1/2((cos A/a)+(cos B/b)+(cos C/c))

myininaya (myininaya):

Use the law of cosines \[a^2=b^2+c^2-2bc \cos(A)\] \[b^2=a^2+c^2-2ac \cos(B)\] \[c^2=a^2+b^2-2ab \cos(C)\]

myininaya (myininaya):

Combine the fractions on the right hand side of the equal sign

myininaya (myininaya):

So you will have \[\frac{\cos(A)}{a}+\frac{\cos(B)}{b}+\frac{\cos(C)}{c}=\frac{bc \cos(A)+ac \cos(B)+ab \cos(C)}{abc}\] and by the way i know who you are @zotactic :) PS

myininaya (myininaya):

So you will also need to multiply that fraction by 2/2 since the other side has a 2 in the bottom

myininaya (myininaya):

\[\frac{2}{2} \cdot \frac{bc \cos(A)+ac \cos(B)+ab \cos(C)}{abc}\]

myininaya (myininaya):

Distribute the 2 on top

myininaya (myininaya):

Now go back to the equations I wrote for you at the beginning and solve each of them for the 2*side*side cos(angle) thingy ma jigger

myininaya (myininaya):

And replace in the new expression I wrote